OpenStudy (anonymous):

An art gallery had 625 visitors in July. In August the gallery had 475 visitors. What is the amount of decrease in the number of visitors from July to August? (Show your work). What is the percent of decrease in the number of visitors from July to August? (Show your work.)

OpenStudy (anonymous):

625 - 475 = 150 625 - 625x = 150 625x = 475 x = 475/625 = 0.76 = 76%

OpenStudy (anonymous):

so is this the the answer for the first question ??

OpenStudy (anonymous):

ohh sorry i get it....

OpenStudy (anonymous):

Wait, scratch that. The answer to the second part is 1 - 0.76, or 0.24, 24%

OpenStudy (anonymous):

what?????????i dont get it..

OpenStudy (anonymous):

475 is 76% of 625, so there was a 24% decrease, since 24% + 76% = 100%

OpenStudy (anonymous):

so the answer for the second part is 24%....

OpenStudy (anonymous):

Yes.